OpenStudy (anonymous):

Solve 2(4 – b) = 2 − b –(b − 6).

OpenStudy (konradzuse):

is it b –(b − 6). or b(b − 6).

OpenStudy (anonymous):

b-(b-6)

OpenStudy (konradzuse):

ok then so distribute 2(4-b) = 8-2b then the 2-b-b-6 is 2-6 = -4 so we getr 8-2b = -4 -8 from both sides -12 = -2b divide -2 on each b = 6

OpenStudy (anonymous):

these are the answers??? b = 0 b = -4 This is an identity and is true for all real numbers. There is no solution.

OpenStudy (konradzuse):

actually I mesed up.

OpenStudy (konradzuse):

2--6 = 8 -2b = 0 b = 0

OpenStudy (konradzuse):

they are = to each other.

OpenStudy (konradzuse):

8-2b = 8-2b